OpenStudy (anonymous):

What is the value of the system determinant for the following system of equations? 6x + y = 4 5x + y - 7 = 0 a)-1 b)1 c)31

OpenStudy (anonymous):

What is the value of the system determinant for the following system of equations? 6x + y = 4 5x + y - 7 = 0 or 5x+y=7. The determinant for this system of equations is 6*1-5*1=1.

OpenStudy (anonymous):

can you show me how to do this?

OpenStudy (anonymous):

@abayomi12 http://openstudy.com/code-of-conduct

OpenStudy (anonymous):